OpenStudy (anonymous):

racheal brought a flute for $265. sales tax was 5.8%.what is the total cost of the flute with tax.

OpenStudy (anonymous):

265/1.058=250.47

OpenStudy (lasttccasey):

Take the 5.8% and turn it into a decimal, .058. Add 1 to get 1.058. Multiple them and theres the total with tax. 265*1.058=$280.73

OpenStudy (anonymous):

tax was 14.5

OpenStudy (lasttccasey):

Tax was $15.37

OpenStudy (lasttccasey):

Only divide 1.058 when tax is already included, \[280.27 / 1.058 = 265\]